OpenStudy (anonymous):

How do you combine terms in (2/x) + (1/x+3) to get (3x+6)/(x^2+3x)?

OpenStudy (phi):

you need a common denominator. one way to get a common denominator is multiply the 2 denominators: x(x+3) now you need to multiply each fraction so you get that denominator. for the first fraction, multiply by (x+3)/(x+3) do you see what to multiply the 2nd fraction by to make its denominator x(x+3) ?

OpenStudy (anonymous):

x/x?

OpenStudy (phi):

yes

OpenStudy (anonymous):

and I can do that because it's just 1 so I'm technically just multiplying by 1?

OpenStudy (phi):

so you get \[ \frac{2}{x}\frac{(x+3)}{(x+3)} + \frac{1}{(x+3)} \frac{x}{x} \] yes, you are multiplying by 1 in each case, now distribute the x in the bottom to get x^2+3x up top, you can simplify

OpenStudy (anonymous):

awesome! That makes so much sense now. Thanks!